The temperature of a cup of cocoa t minutes after it is made is given by the function C(t) = (a + b)(0.95)^t, where a is the temperature of the room, in degrees Celsius, and a + b is the initial temperature of the cocoa. in degrees Celsius. If the cocoa is initially 80°C and the temperature of the room is 20°C, what will the temperature of the cocoa be 2 minutes later, rounded to the nearest degree Celsius?
Answer by ikleyn(52756)

.
The temperature of a cup of cocoa t minutes after it is made is given by the function C(t) = (a + b)(0.95)^t,
where a is the temperature of the room, in degrees Celsius, and a + b is the initial temperature of the cocoa.
in degrees Celsius. If the cocoa is initially 80°C and the temperature of the room is 20°C,
what will the temperature of the cocoa be 2 minutes later, rounded to the nearest degree Celsius?
~~~~~~~~~~~~~~~~~


The formula for the temperature  C(t)  is written incorrectly in your post.


To see many similar solved problems of this type,  look into the lesson
    - Solving problem on Newton Law of cooling
in this site.

Learn the subject from there.


Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Logarithms".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.


===================


Comment from student: 𝐶(0)=𝑎+𝑏=80°𝐶 𝑎=20°𝐶 𝑏=80°𝐶−20°𝐶 𝑏=60°𝐶 𝐶(𝑡)=20+60(0.95)^𝑡 𝐶(2)=20+60(0.95)^2 𝐶(2)=20+60[0.9025] 𝐶(2)=20+54.15 𝐶(2)=74.15°𝐶≈74°𝐶



My response :   Hey,  I do not need your solution,  so you posted it to me to not purpose.

You better learn to write the formulas correctly . . .